The price of a perpetuity with monthly cash flows of $50 when the current interest rate is 7.24% p.a. compounding monthly and the first cash flow occurs at the end of month #2 is:
Answer #1

Perpetuity is periodic cash inflow of even amount for indefinite period of time, in this question we need to find the PV of the cash inflows.

--------------------------------------------------------------------------------------------------------------------------

Interest rate per month = 7.24/12 = .6033% per month

Price of the perpetuity = Periodic payment/ interest rate

                                           = 50/0.006033

                                            = $8287.75

Since first payment is received at the end of month 2, we need to discount further to the PV

So price of the perpetuity = 8287.75/(1 + 0.006033)

                                                = 8287.75/1.006033

                                                = $8238.05

--------------------------------------------------------------------------------------------------------------------------

Hope that helps.

Feel free to comment if you need further assistance J